In a vacuum, two particles have charges of q1 and q2, where q1 = +5.0C. They are separated by a distance of 0.24 m, and particle 1 experiences an attractive force of 3.1 N. What is the value of q2, with its sign?

Thanks for the help!
Answer #1

Here ,

as the force is attractive , the value of q2 is negative

hence , for the force

-3.1 = k * q1 * q2/d^2

-3.1 = 9 *10^9 * 5 * q2/0.24^2

solving for q2

q2 = -3.97 *10^-12 C

the charge q2 is -3.97 *10-12 C